OpenStudy (anonymous):

simplify (4x)^-3 (4x)^-3=____

jhonyy9 (jhonyy9):

how do you think it ?

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

to simplifie this ...

jhonyy9 (jhonyy9):

so if you know that a^(-1) = 1/a so than what you will get in this case ?

OpenStudy (anonymous):

is it 64?

jhonyy9 (jhonyy9):

sure ?

OpenStudy (anonymous):

do you know what the answer is?

jhonyy9 (jhonyy9):

i know it but you need to solve it and i like help you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 64x

jhonyy9 (jhonyy9):

why ?

jhonyy9 (jhonyy9):

so you know that a^(-1) = 1/a

jhonyy9 (jhonyy9):

so than (4x)^(-3) = 1/(4x)^3 yes ?

jhonyy9 (jhonyy9):

and you know again that (4x)^3 = 4^3 *x^3 ok ?

jhonyy9 (jhonyy9):

4^3 = ?

jhonyy9 (jhonyy9):

so than (4x)^(-3) = 1/(4x)^3 = 1/(64x^3) ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer is 64^3?